OpenStudy (anonymous):

simplify the expression (1-3x+5)+(7x+6-3x)

OpenStudy (anonymous):

ok let raequelle get some practice :)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

it is x+12 first bracket resolves to 6-3x second resolves to 4x+6 (6-3x) + (4x+6) = -3x+4x+6+6 = x+12